OpenStudy (anonymous):

-3x^2 + 52x + 378 = 0 ... slove ths by completing the square method.

ganeshie8 (ganeshie8):

start by factoring out the GCF

OpenStudy (anonymous):

??? opss sry, it must be sloved with spilitting the middle term , ......

ganeshie8 (ganeshie8):

yes, thats the second step

ganeshie8 (ganeshie8):

first step for all these methods is to factor out GCF if any..

ganeshie8 (ganeshie8):

however there is no GCF here :) so lets continue wid splitting the middle term..

ganeshie8 (ganeshie8):

before doing that, just make sure u dont have any typo in the middle term is it really 52, and not 42 ? ??

OpenStudy (anonymous):

let me check , a sec plz

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

yeah! i got 52 only, .... actually question is , a vehicle travels at a certain average speed for a distance of 63km and thn travels a ditance of 72km at an average speed of 6km/h more than its original speed. if it takes 3 hours to complete the total journey, what si its original averge speed???

OpenStudy (anonymous):

ths is the actually question, after slovign i got tht above question answer, .......

ganeshie8 (ganeshie8):

one sec..

OpenStudy (anonymous):

hehe maybe i sloved it wrong, ... coudl u plz slove it

OpenStudy (anonymous):

sure...

ganeshie8 (ganeshie8):

\(\large \mathbb{ \frac{63}{x} + \frac{72}{x+6} = 3 }\)

ganeshie8 (ganeshie8):

is that the equation u got ?

OpenStudy (anonymous):

yeah! yes i got tht 1 only, .....

ganeshie8 (ganeshie8):

okay lets simplify this and solve :)

OpenStudy (anonymous):

sure ...

ganeshie8 (ganeshie8):

\(\large \mathbb{ \frac{63}{x} + \frac{72}{x+6} = 3 }\) \(\large \mathbb{ \frac{63(x+6)}{x(x+6)} + \frac{72x}{x(x+6)} = 3 }\) \(\large \mathbb{ 63(x+6)+ 72x = 3x(x+6) }\)

ganeshie8 (ganeshie8):

fine ?

OpenStudy (anonymous):

a sec let me check, ....... had anet probelm just now

OpenStudy (anonymous):

accha!!! so we must multiply x(x+6) to 3 also,!!! didn't knew :P

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

yeah! wt after tht???

ganeshie8 (ganeshie8):

\(\large \mathbb{ \frac{63}{x} + \frac{72}{x+6} = 3 }\) \(\large \mathbb{ \frac{63(x+6)}{x(x+6)} + \frac{72x}{x(x+6)} = 3 }\) \(\large \mathbb{ 63(x+6)+ 72x = 3x(x+6) }\) \(\large \mathbb{ 21(x+6)+ 24x = x(x+6) }\) \(\large \mathbb{ 21x+126+ 24x = x^2+6x }\) \(\large \mathbb{ x^2-39x -126 = 0 }\)

ganeshie8 (ganeshie8):

you wud get that quadratic... usko sovle kar

ganeshie8 (ganeshie8):

if u want, u may split the mid term : -126 = -42*3 -39 = -42+3

ganeshie8 (ganeshie8):

\(\large \mathbb{ x^2-39x -126 = 0 }\) \(\large \mathbb{ (x+3)(x-42) = 0 }\)

ganeshie8 (ganeshie8):

\(\large \mathbb{ x=-3~~~ x = 42 }\)

ganeshie8 (ganeshie8):

discard the negative value so the original average speed during the first 63km is \(42\)

ganeshie8 (ganeshie8):

\(42 \) kmph

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

oopsss a doubt , .... after tht >>> 21(x+6)24x= x (x+6) 21x+126+24x= x^2 +6x... thr above hw did u / y did u add 24x again?

OpenStudy (anonymous):

anyways thnks

ganeshie8 (ganeshie8):

hey u still have that doubt ?

OpenStudy (anonymous):

yeah! :( , poor in maths

ganeshie8 (ganeshie8):

no you're very good, and thats a good question to ask actually.. let me give it to u bit detailed

ganeshie8 (ganeshie8):

\(\large \mathbb{ 63(x+6)+ 72x = 3x(x+6) }\) divide both sides by 3 : wat do u get ?

OpenStudy (anonymous):

omg!!! sm1 prasied me tht alos in maths , ....... let me slove wait, .... 21(x+6) + 24x = x(x+6) , ...

OpenStudy (anonymous):

huraay! i got it , ..... by myself

ganeshie8 (ganeshie8):

yes thats right ! you're smart lol

OpenStudy (anonymous):

:P smart , heheh applying butter on me, ... anyways wt abt ths question>

ganeshie8 (ganeshie8):

wat question ha

ganeshie8 (ganeshie8):

u already solved it by urself u got 24x right -_-

OpenStudy (anonymous):

PT and QT are two tangents to the circle, if angle PTQ = 50 degree, find angle PSQ and angle OPQ. ....

ganeshie8 (ganeshie8):

oh another question ok go ahead..

OpenStudy (anonymous):

another quetion men, ....... tht question is already slove :P

ganeshie8 (ganeshie8):

ok ok just realized lol :D

ganeshie8 (ganeshie8):

draw it out

OpenStudy (anonymous):

:P , ... hw to slove tht 1

ganeshie8 (ganeshie8):

start by drawing the circle and tangents and also label them

OpenStudy (anonymous):

sec plz

ganeshie8 (ganeshie8):



OpenStudy (anonymous):